# Signed Angle Rigid Graphs for Network Localization and Formation Control

### v1

@@ Abstract
Graph rigidity theory studies the capability of a graph embedded in the Euclidean space to constrain its global geometric shape via local constraints among nodes and edges, and has been widely exploited in network localization and formation control. In recent years, the traditional rigidity theory has been extended by considering new types of local constraints such as bearing, angle, ratio of distance, etc. Among them, the signed angle constraint received extensive attention since it is practically measurable and independent of the global coordinate frame. However, existing studies on signed angle rigidity always consider special graph structures, which are actually not necessary. This paper presents a comprehensive combinatorial analysis in terms of graphs and angle index sets for signed angle rigidity. We show that Laman graphs equivalently characterize minimally signed angle rigid graphs. Moreover, we propose a method to construct the minimal set of signed angle constraints in a Laman graph to effectively ensure signed angle rigidity. These results are finally applied to distributed network localization and formation stabilization problems, respectively, where each agent only has access to signed angle measurements.
@@ I Introduction
The traditional rigidity theory (distance rigidity theory) studies the characterizability of a geometric shape by distance constraints among neighboring vertices in the graph, and has been widely employed in network localization and formation control [1, 3, 2, 4, 5, 6, 7]. Inspired by this, rigidity theories based on other types of constraints (such as bearings [8, 9], angles [10, 11], ratio of distance [14], etc.) have been further developed to study network localization [16, 19, 15, 17, 18] and formation stabilization [8, 9, 10, 11, 12, 13] via different sensing measurements. A geometric shape specified by different types of constraints may have different degrees of freedom and require different graphical conditions to ensure a unique shape [22].
@@
In distance-based formation control, infinitesimal distance rigidity has been adopted most widely since it has an equivalent rank condition on the rigidity matrix [23], which plays an essential role in the analysis of the gradient-based formation controller [24, 3, 25, 26]. However, infinitesimal distance rigidity does not guarantee the shape to be uniquely determined by distance constraints [27], which usually leads to local convergence of the formation. On the other hand, although global distance rigidity implies a unique shape, it usually requires a large number of edges, and the corresponding formation controller to achieve global convergence is yet to be developed [28]. In [9], infinitesimal bearing rigidity was shown to share the same graphical conditions as infinitesimal distance rigidity (i.e., Laman graphs [29, 27]), while guaranteeing that the geometric shape is uniquely determined by bearings. Guided by this observation, the authors in [9, 15] proposed novel distributed protocols for bearing-based formation stabilization and sensor network localization. However, bearing constraints are associated with the global coordinate frame, which limits its applicability.
@@
Recently, angle rigidity theory has garnered significant interest, as angle constraints are independent of the coordinate frame [16, 21, 20, 17, 18], and an angle-constrained formation exhibits higher degrees of freedom than distance and bearing-based approaches [10, 11]. In [10], the authors established the equivalence between infinitesimal angle rigidity, infinitesimal bearing rigidity, and infinitesimal distance rigidity in the plane. However, it is unclear whether the shape of an infinitesimally angle rigid framework can be uniquely determined by angles. In [11], the authors reformulated angle rigidity theory by considering signed angle constraints. However, signed angle rigidity was studied on angularities with predefined sets of angle constraints. Such a setting benefits for excluding redundant angles; however, it hinders the connection between the graph structure and the rigidity property.
@@
In this paper, we study the conditions on graphs and the minimal set of signed angle constraints required to ensure angle rigidity and the uniqueness of the shape. Similar issues related to angle rigidity theory have been investigated in the literature. In [30, 31], the authors explored the combinatorial conditions of angle rigidity. In [18, 11, 10], the authors reported a class of particular graphs which ensure a unique shape under angle constraints. However, these works fail to characterize the equivalent graphical conditions for angle rigidity and the minimal set of angle constraints to guarantee a unique shape. Considering the large body of existing works in bearing rigidity theory, we believe that establishing the relationship between signed angle rigidity and bearing rigidity would be a worthwhile exploration.
@@
Our contributions can be summarized as follows. (i). Enlightened by angle rigidity theory in [10, 11], we develop the signed angle rigidity theory in the context of frameworks, where all signed angles within the framework are utilized. Under this setting, it is shown that (infinitesimal, global) signed angle rigidity and (infinitesimal, global) bearing rigidity are completely equivalent, which implies that Laman graphs are exactly minimally signed angle rigid graphs; see Section IIIII. (ii). Similarly to the bearing case, we prove that infinitesimal signed angle rigidity is necessary and sufficient to determine a unique shape; see Theorem 3. (iii). We provide a combinatorial analysis on the set of angle constraints and develop Algorithm 1 to generate a minimal angle index set (AIS) to characterize the rigidity property of a framework; see Section IV. (iv). Based on the theory developed, we solve the problems of network localization and formation stabilization, respectively, where each agent only measures local signed angle measurements associated with its neighbors. Two distributed algorithms with global convergence guarantee are proposed, respectively, where the conditions on the sensing graph are much milder than those in [18, 10, 17, 11]; see Section VVI.
@@ III Graphs for Signed Angle Rigidity
In this section, we establish the equivalence between signed angle constraints and bearing constraints in the sense that a framework $(\mathcal{G},p)$ is IBR, GBR, and BR if and only if it is ISAR, GSAR, and SAR, respectively, under which we derive the necessary and sufficient conditions for frameworks to be determined by signed angles only, and graphs to be signed angle rigidity.
@@ III-B Equivalence on GSAR and GBR
Combining Theorems 1 and 2, together with [9, Theorem 3] and [9, Theorem 5], the following results hold.
@@ III-D Signed Angle Rigid Graphs
In [29, Lemma 2], IBR has been shown as a generic property of graphs. More specifically, if $(\mathcal{G},p)$ is IBR for a configuration $p\in\mathbb{R}^{2n}$, then $(\mathcal{G},q)$ is IBR for almost all $q\in\mathbb{R}^{2n}$. According to [9], such a graphic property holds for GBR and BR as well. From the relationship between bearing rigidity and signed angle rigidity shown in Fig. 3, it can be observed that ISAR, GSAR, and SAR are essentially determined by the graph rather than the configuration, thus are generic properties of graphs. We summarize this result in the following theorem.
@@
Combining Theorem 1 and [29, Theorem 2], we present the following result, which implies the shape of the graphs containing Laman spanning subgraphs can be determined by signed angle constraints for almost all configurations.
@@
For any ISAR framework $(\mathcal{G},p)$ in $\mathbb{R}^{2}$, we can find an ISAR Laman spanning subframework in $(\mathcal{G},p)$ based on the distance rigidity matrix [27], where a framework is said to be the Laman framework if it is embedded by a Laman graph. Specifically, considering an ISAR framework $(\mathcal{G},p)$ in $\mathbb{R}^{2}$, according to [9, Theorem 8] and Theorem 1, $(\mathcal{G},p)$ is infinitesimally distance rigid. Then there are $32n-$ 3 linearly independent rows in the distance rigidity matrix. According to [27, Lemma 2.3], the edges corresponding to these $32n-$ 3 rows are also the edges of an ISAR Laman spanning subframework.
@@ IV-A Minimal GAISs for Signed Angle Rigidity
According to Theorem 3 and Lemma 5, we know that for an ISAR framework $(\mathcal{G},p)$ in $\mathbb{R}^{2}$, $\bar{\mathcal{T}}_{\mathcal{G}}\subseteq\mathcal{T}_{\mathcal{G}}$ is a GAIS if and only if $\mathscr{E}_{\bar{\mathcal{T}}_{\mathcal{G}}}(p)=\mathscr{E}(p)$. Since $p$ is non-degenerate, $\mathscr{E}_{\bar{\mathcal{T}}_{\mathcal{G}}}(p)$ is a 4-dimensional smooth manifold. According to [33, Proposition 3.10], the tangent space of $\mathscr{E}_{\bar{\mathcal{T}}_{\mathcal{G}}}(p)$ at $p$ is 4-dimensional, which indicates that $\operatorname{rank}(R^{\bar{\mathcal{T}}_{\mathcal{G}}}_{S}(p))=2n-4$. As a result, one has $|\bar{\mathcal{T}}_{\mathcal{G}}|\geq 2n-4$.
@@
On the one hand, according to the proof of Theorem 7, we observe that for a framework $(\mathcal{G},p)$ in $\mathbb{R}^{2}$, the breadth first search method (BFS) [36, Section 20.2] can be applied to extract a spanning tree from the angle index graph $(\mathcal{V}_{A}(\mathcal{E}),\mathcal{E}_{A}(\mathcal{T}_{\mathcal{G}}))$. The resulting AIS $\bar{\mathcal{T}}_{\mathcal{G}}$ associated with this spanning tree is angle connected and satisfy $|\bar{\mathcal{T}}_{\mathcal{G}}|=|\mathcal{E}|-1$. On the other hand, according to Theorem 3-5, a framework can be uniquely determined by signed angles if and only if it contains an ISAR Laman subframework. Based on these observations, we present Algorithm 1 to construct a minimal GAIS $\bar{\mathcal{T}}_{\mathcal{G}}$ for arbitrary ISAR frameworks, which enables unique shape while satisfying $|\bar{\mathcal{T}}_{\mathcal{G}}|=2n-4$.
@@
2:  Find a spanning tree $(\mathcal{V}^{tree}_{A},\mathcal{E}^{tree}_{A})$ within the angle index graph $(\mathcal{V}_{A}(\mathcal{L}_{n}),\mathcal{E}_{A}(\mathcal{T}_{\mathcal{L}_{n}}))$ by BFS [36, Section 20.2].
@@
Algorithm 1 is a polynomial-time algorithm. As noted in Remark 3, Line 1 can be achieved via Gaussian elimination [37, Section 3.2], with a time complexity of $\mathcal{O}(|\mathcal{E}|n^{2})$ when $m\geq 2n$, and $\mathcal{O}(|\mathcal{E}|^{2}n)$ otherwise. Since the worst-case time complexity of Lines 2–5 is $\mathcal{O}(n^{2})$ (when $(\mathcal{V}_{A}(\mathcal{L}_{n}),\mathcal{E}_{A}(\mathcal{T}_{\mathcal{L}_{n}}))$ is a complete graph) and $|\mathcal{E}|\leq n(n-1)/2$, the worst-case time complexity of Algorithm 1 is $\mathcal{O}(n^{5})$.
@@ V-A Problem Formulation and Network Localizability
V-B A Distributed Algorithm for Network Localization
@@ VI-A The Control Law and Stability Analysis
It can be observed that (36) is the same as the consensus algorithm in [39, Lemma 1.3] under a connected undirected graph $\mathcal{G}$. Therefore, $\beta_{ij},(i,j)\in\mathcal{E}$ converges to zero exponentially fast for any initial points $\beta_{i}(0)\in[0,2\pi),i\in\mathcal{V}$. That is, the body frames of agents will achieve coordinate alignment exponentially. Using similar techniques in the proof of [9, Theorem 15], it can be concluded that the controller (34) will drive $b_{ij}(p),(i,j)\in\mathcal{E}$ to either $\mathscr{R}_{o}(\theta)b_{ij}(p^{*})$ (for almost all initial values) or $-\mathscr{R}_{o}(\theta)b_{ij}(p^{*})$ (for the remaining case) exponentially fast, where $\theta\in[0,2\pi)$. It follows that $\alpha_{ijk}(p)\rightarrow\alpha_{ijk}(p^{*}),(i,j,k)\in\mathcal{T}_{\mathcal{G}}$ as $t\rightarrow\infty$ from any initial values. According to Theorem 3, the position $p(t)$ of the formation converges to the manifold $\mathscr{E}(p^{*})$ globally. ∎
@@ VIII-B Proof of Theorem 1
According to [10, Theorem 5], we know that for a framework $(\mathcal{G},p)$ in $\mathbb{R}^{2}$, $\operatorname{rank}(R_{B}(p))=2n-3$ if and only if $\operatorname{rank}(R^{\mathcal{T}_{\mathcal{G}}}_{A}(p))=2n-4$, where $R^{\mathcal{T}_{\mathcal{G}}}_{A}(p)$ is the rigidity matrix defined in [10] using unsigned angles (i.e., the cosine of the angles) as the entries of the rigidity function [^1] In this paper, we use the triple $(i,j,k)$ to index the unsigned angle between $b_{ji}$ and $b_{jk}$, which is consistent with the definition in [10].. Therefore, it suffices to show that $\operatorname{rank}(R^{\mathcal{T}_{\mathcal{G}}}_{A}(p))=2n-4$ if and only if $\operatorname{rank}(R^{\mathcal{T}_{\mathcal{G}}}_{S}(p))=2n-4$.
From Lemma 8 and [10], we know that if $(\mathcal{G},p)$ satisfies $\operatorname{rank}(R^{\mathcal{T}_{\mathcal{G}}}_{A}(p))=2n-4$ or $\operatorname{rank}(R^{\mathcal{T}_{\mathcal{G}}}_{S}(p))=2n-4$, then for any vertex $j\in\mathcal{V}$, there exist at least two distinct neighbors $i,k\in\mathcal{N}_{j}$ such that $|b^{\top}_{ji}b_{jk}|\neq 1$. As a result, when the framework satisfies $\operatorname{rank}(R^{\mathcal{T}_{\mathcal{G}}}_{A}(p))=2n-4$ or $\operatorname{rank}(R^{\mathcal{T}_{\mathcal{G}}}_{S}(p))=2n-4$, we can find an angle connected set $\mathcal{T}^{*}_{\mathcal{G}}\subseteq\mathcal{T}_{\mathcal{G}}$ by [10, Algorithm 1], which satisfies $\sin\alpha_{ijk}\neq 0$ for $(i,j,k)\in\mathcal{T}^{*}_{\mathcal{G}}$ and $\operatorname{rank}(R^{\mathcal{T}^{*}_{\mathcal{G}}}_{A}(p))=\operatorname{rank}(R^{\mathcal{T}_{\mathcal{G}}}_{A}(p))$. Moreover, with virtue of Lemma 7 (will be proposed later with its proof independence of this theorem), one has $\operatorname{rank}(R^{\mathcal{T}^{*}_{\mathcal{G}}}_{S}(p))=\operatorname{rank}(R^{\mathcal{T}_{\mathcal{G}}}_{S}(p))$.

### v2

@@ Abstract
Graph rigidity theory studies the capability of a graph embedded in the Euclidean space to constrain its global geometric shape via local constraints among nodes and edges, and has been widely exploited in network localization and formation control. In recent years, the traditional rigidity theory has been extended by considering new types of local constraints such as bearing, angle, ratio of distance, etc. Among them, the signed angle constraint has received extensive attention, since it is practically measurable and independent of the global coordinate frame. However, the relevant studies always consider special graph structures, which are sufficient but not necessary for signed angle rigidity. This paper presents a comprehensive combinatorial analysis in terms of graphs and angle index sets for signed angle rigidity. We show that Laman graphs equivalently characterize minimally signed angle rigid graphs. Moreover, we propose a method to construct the minimal set of signed angle constraints in a Laman graph to effectively ensure signed angle rigidity. These results are finally applied to distributed network localization and formation stabilization problems, respectively, where each agent only has access to signed angle measurements.
@@ I Introduction
The traditional graph rigidity theory studies the characterizability of a geometric shape by distance constraints among neighboring vertices in the graph, and has been widely employed in multi-agent coordination problems such as network localization [1, 2, 3] and formation control [4, 5, 6, 7, 8]. Inspired by this, rigidity theories based on other types of constraints (such as bearings [9, 10], angles [11, 12], ratio of distance [13], etc.) have been further developed to achieve network localization [14, 15, 16, 17, 18] and formation stabilization [9, 10, 11, 12, 19, 20, 13] via different sensing measurements. A geometric shape specified by different types of constraints may require different graphical conditions to ensure a unique shape [21]. As a result, different conditions on the network topology have been proposed depending on the type of measurements utilized.
@@
In distance-based multi-agent coordination problems, infinitesimal distance rigidity has been adopted most widely since it has an equivalent rank condition on the rigidity matrix [22], which plays an essential role in the analysis of the gradient-based controller [23, 5, 24, 25]. However, infinitesimal distance rigidity does not guarantee the shape to be uniquely determined by distance constraints [26], which usually leads to local convergence of the controller. On the other hand, although global distance rigidity implies a unique shape, and has been favored in sensor network localization [1], it usually requires a large number of edges. Moreover, the formation controller achieving global convergence under globally rigid graphs is yet to be designed [27]. In [10], infinitesimal bearing rigidity was shown to share the same graphical conditions with infinitesimal distance rigidity (i.e., Laman graphs [28, 26]), while guaranteeing that the geometric shape is uniquely determined by bearings. Guided by this observation, the authors in [10, 16] proposed novel distributed protocols for bearing-based formation stabilization and sensor network localization. However, bearing constraints are associated with the global coordinate frame, which limits its applicability.
@@
Recently, angle rigidity theory has garnered significant interest, as angle measurements are independent of the coordinate frame [14, 29, 30, 17, 18, 31], and an angle-constrained formation exhibits higher degrees of freedom than distance and bearing-based approaches [11, 12]. In [11], the authors established the equivalence between infinitesimal angle rigidity, infinitesimal bearing rigidity, and infinitesimal distance rigidity in the plane. However, it is unclear whether the shape of an infinitesimally angle rigid framework can be uniquely determined by angles. In [12], the authors reformulated angle rigidity theory by considering signed angle constraints. Unfortunately, the signed angle rigidity property was studied on angularities with predefined sets of angle constraints. Such a setting benefits for excluding redundant angles, but in the meanwhile, hinders the connection between the graph structure and the signed angle rigidity property.
@@
In this paper, we study the conditions on graphs and the minimal set of signed angle constraints required to ensure signed angle rigidity and uniqueness of the shape. Similar issues related to angle rigidity theory have been investigated in the literature. In [32, 33], the authors explored the combinatorial conditions of angle rigidity. In [18, 12, 11], the authors reported a class of particular graphs which ensure a unique shape under angle constraints. Nevertheless, these works fail to characterize the equivalent graphical conditions for angle rigidity and the minimal set of angle constraints to guarantee a unique shape. Considering the large body of existing works in bearing rigidity theory, we believe that establishing the relationship between signed angle rigidity and bearing rigidity would be a worthwhile exploration.
@@
Our contributions can be summarized as follows. (i). Enlightened by angle rigidity theory in [11, 12], we develop the signed angle rigidity theory in the context of frameworks, where all signed angles within the framework are utilized. Under this setting, it is shown that (infinitesimal, global) signed angle rigidity and (infinitesimal, global) bearing rigidity are completely equivalent, which implies that Laman graphs are exactly minimally signed angle rigid graphs; see Section IIIII. (ii). Similarly to the bearing case, we prove that infinitesimal signed angle rigidity is sufficient and necessary for a framework to determine its shape uniquely; see Theorem 3. (iii). We provide a combinatorial analysis on the set of angle constraints and develop Algorithm 1 to generate a minimal angle index set (AIS) to characterize the rigidity property of a framework; see Section IV. (iv). Based on the theory developed, we solve the problems of network localization and formation stabilization, respectively, where each agent only has access to local signed angle measurements associated with its neighbors. Two distributed algorithms with global convergence guarantee are proposed, respectively, where the conditions on the sensing graph are much milder than those in [18, 11, 17, 12]; see Section VVI.
@@ III Graphs for Signed Angle Rigidity
In this section, we establish the equivalence between signed angle constraints and bearing constraints in the sense that a framework $(\mathcal{G},p)$ is IBR, GBR, and BR if and only if it is ISAR, GSAR, and SAR, respectively. Based on these relationships, we derive necessary and sufficient graphical conditions for signed angle rigidity.
@@ III-B Equivalence on GSAR and GBR
Combining Theorems 1 and 2, together with [10, Theorem 3] and [10, Theorem 5], the following results hold.
@@ III-D Signed Angle Rigid Graphs
In [28, Lemma 2], IBR has been shown as a generic property of graphs. More specifically, if $(\mathcal{G},p)$ is IBR for a configuration $p\in\mathbb{R}^{2n}$, then $(\mathcal{G},q)$ is IBR for almost all $q\in\mathbb{R}^{2n}$. According to [10], such a graphic property holds for GBR and BR as well. From the relationship between bearing rigidity and signed angle rigidity shown in Fig. 3, it can be observed that ISAR, GSAR, and SAR are essentially determined by the graph rather than the configuration, thus are generic properties of graphs. We summarize this result in the following theorem.
@@
Combining Theorem 1 and [28, Theorem 2], we present the following result, which implies the shape of the graphs containing Laman spanning subgraphs can be determined by signed angle constraints for almost all configurations.
@@
For any ISAR framework $(\mathcal{G},p)$ in $\mathbb{R}^{2}$, we can find an ISAR Laman spanning subframework in $(\mathcal{G},p)$ based on the distance rigidity matrix [26], where a framework is said to be the Laman framework if it is embedded by a Laman graph. Specifically, considering an ISAR framework $(\mathcal{G},p)$ in $\mathbb{R}^{2}$, according to [10, Theorem 8] and Theorem 1, $(\mathcal{G},p)$ is infinitesimally distance rigid. Then there are $32n-$ 3 linearly independent rows in the distance rigidity matrix. According to [26, Lemma 2.3], the edges corresponding to these $32n-$ 3 rows are also the edges of an ISAR Laman spanning subframework.
@@ IV-A Minimal GAISs for Signed Angle Rigidity
According to Theorem 3 and Lemma 5, we know that for an ISAR framework $(\mathcal{G},p)$ in $\mathbb{R}^{2}$, $\bar{\mathcal{T}}_{\mathcal{G}}\subseteq\mathcal{T}_{\mathcal{G}}$ is a GAIS if and only if $\mathscr{E}_{\bar{\mathcal{T}}_{\mathcal{G}}}(p)=\mathscr{E}(p)$. Since $p$ is non-degenerate, $\mathscr{E}_{\bar{\mathcal{T}}_{\mathcal{G}}}(p)$ is a 4-dimensional smooth manifold. According to [36, Proposition 3.10], the tangent space of $\mathscr{E}_{\bar{\mathcal{T}}_{\mathcal{G}}}(p)$ at $p$ is 4-dimensional, which indicates that $\operatorname{rank}(R^{\bar{\mathcal{T}}_{\mathcal{G}}}_{S}(p))=2n-4$. As a result, one has $|\bar{\mathcal{T}}_{\mathcal{G}}|\geq 2n-4$.
@@
On the one hand, according to the proof of Theorem 7, we observe that for a framework $(\mathcal{G},p)$ in $\mathbb{R}^{2}$, the breadth first search method (BFS) [37, Section 20.2] can be applied to extract a spanning tree from the angle index graph $(\mathcal{V}_{A}(\mathcal{E}),\mathcal{E}_{A}(\mathcal{T}_{\mathcal{G}}))$. The resulting AIS $\bar{\mathcal{T}}_{\mathcal{G}}$ associated with this spanning tree is angle connected and satisfy $|\bar{\mathcal{T}}_{\mathcal{G}}|=|\mathcal{E}|-1$. On the other hand, according to Theorem 3-5, a framework can be uniquely determined by signed angles if and only if it contains an ISAR Laman subframework. Based on these observations, we present Algorithm 1 to construct a minimal GAIS $\bar{\mathcal{T}}_{\mathcal{G}}$ for arbitrary ISAR frameworks, which enables unique shape while satisfying $|\bar{\mathcal{T}}_{\mathcal{G}}|=2n-4$.
@@
2:  Find a spanning tree $(\mathcal{V}^{tree}_{A},\mathcal{E}^{tree}_{A})$ within the angle index graph $(\mathcal{V}_{A}(\mathcal{L}_{n}),\mathcal{E}_{A}(\mathcal{T}_{\mathcal{L}_{n}}))$ by BFS [37, Section 20.2].
@@
Algorithm 1 is a polynomial-time algorithm. As noted in Remark 3, Line 1 can be achieved via Gaussian elimination [38, Section 3.2], with a time complexity of $\mathcal{O}(|\mathcal{E}|n^{2})$ when $m\geq 2n$, and $\mathcal{O}(|\mathcal{E}|^{2}n)$ otherwise. Since the worst-case time complexity of Lines 2–5 is $\mathcal{O}(n^{2})$ (when $(\mathcal{V}_{A}(\mathcal{L}_{n}),\mathcal{E}_{A}(\mathcal{T}_{\mathcal{L}_{n}}))$ is a complete graph) and $|\mathcal{E}|\leq n(n-1)/2$, the worst-case time complexity of Algorithm 1 is $\mathcal{O}(n^{5})$.
@@ V-A Problem Formulation and Network Localizability
V-B A Distributed Localization Algorithm
@@ VI-A The Control Law and Stability Analysis
It can be observed that (36) is the same as the consensus algorithm in [40, Lemma 1.3] under a connected undirected graph $\mathcal{G}$. Therefore, $\beta_{ij},(i,j)\in\mathcal{E}$ converges to zero exponentially fast for any initial points $\beta_{i}(0)\in[0,2\pi),i\in\mathcal{V}$. That is, the body frames of agents will achieve coordinate alignment exponentially. Using similar techniques in the proof of [10, Theorem 15], it can be concluded that the controller (34) will drive $b_{ij}(p),(i,j)\in\mathcal{E}$ to either $\mathscr{R}_{o}(\theta)b_{ij}(p^{*})$ (for almost all initial values) or $-\mathscr{R}_{o}(\theta)b_{ij}(p^{*})$ (for the remaining case) exponentially fast, where $\theta\in[0,2\pi)$. It follows that $\alpha_{ijk}(p)\rightarrow\alpha_{ijk}(p^{*}),(i,j,k)\in\mathcal{T}_{\mathcal{G}}$ as $t\rightarrow\infty$ from any initial values. According to Theorem 3, the position $p(t)$ of the formation converges to the manifold $\mathscr{E}(p^{*})$ globally. ∎
@@ VIII-B Proof of Theorem 1
According to [11, Theorem 5], we know that for a framework $(\mathcal{G},p)$ in $\mathbb{R}^{2}$, $\operatorname{rank}(R_{B}(p))=2n-3$ if and only if $\operatorname{rank}(R^{\mathcal{T}_{\mathcal{G}}}_{A}(p))=2n-4$, where $R^{\mathcal{T}_{\mathcal{G}}}_{A}(p)$ is the rigidity matrix defined in [11] using unsigned angles (i.e., the cosine of the angles) as the entries of the rigidity function [^1] In this paper, we use the triple $(i,j,k)$ to index the unsigned angle between $b_{ji}$ and $b_{jk}$, which is consistent with the definition in [11].. Therefore, it suffices to show that $\operatorname{rank}(R^{\mathcal{T}_{\mathcal{G}}}_{A}(p))=2n-4$ if and only if $\operatorname{rank}(R^{\mathcal{T}_{\mathcal{G}}}_{S}(p))=2n-4$.
From Lemma 8 and [11], we know that if $(\mathcal{G},p)$ satisfies $\operatorname{rank}(R^{\mathcal{T}_{\mathcal{G}}}_{A}(p))=2n-4$ or $\operatorname{rank}(R^{\mathcal{T}_{\mathcal{G}}}_{S}(p))=2n-4$, then for any vertex $j\in\mathcal{V}$, there exist at least two distinct neighbors $i,k\in\mathcal{N}_{j}$ such that $|b^{\top}_{ji}b_{jk}|\neq 1$. As a result, when the framework satisfies $\operatorname{rank}(R^{\mathcal{T}_{\mathcal{G}}}_{A}(p))=2n-4$ or $\operatorname{rank}(R^{\mathcal{T}_{\mathcal{G}}}_{S}(p))=2n-4$, we can find an angle connected set $\mathcal{T}^{*}_{\mathcal{G}}\subseteq\mathcal{T}_{\mathcal{G}}$ by [11, Algorithm 1], which satisfies $\sin\alpha_{ijk}\neq 0$ for $(i,j,k)\in\mathcal{T}^{*}_{\mathcal{G}}$ and $\operatorname{rank}(R^{\mathcal{T}^{*}_{\mathcal{G}}}_{A}(p))=\operatorname{rank}(R^{\mathcal{T}_{\mathcal{G}}}_{A}(p))$. Moreover, with virtue of Lemma 7 (will be proposed later with its proof independence of this theorem), one has $\operatorname{rank}(R^{\mathcal{T}^{*}_{\mathcal{G}}}_{S}(p))=\operatorname{rank}(R^{\mathcal{T}_{\mathcal{G}}}_{S}(p))$.